if a letter is chosen at random from the 10 letters of the word statistics wat is probability that it is a vowel?


1
Expert's answer
2021-05-07T10:31:01-0400

Given, "n=10" , Total number of vowels in word ' STATISTICS'=3


Let p denote the probability of vowel "p=\dfrac{5}{26}=0.192"

Then "q=1-p=1-0.192=0.8076"


Probability that selected letter is vowel-

"=^{10}C_3 p^3q^{10-3}\\=^{10}C_3(0.192)^3(0.8076)^7\\=120\times0.00707\times 0.224\\=0.19"